OpenStudy (anonymous):

Use the half angle formula to solve (sin(4x))^4

OpenStudy (anonymous):

No equals sign --> nothing to solve.

OpenStudy (anonymous):

But here's some formulas if you'd like to make some substitutions. http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

Use the half angle formula to solve (Sin(4x))^2=________-1/2cos(_________x)+1/8cos(16x)

OpenStudy (anonymous):

Its basically to fill in the blank in the question

OpenStudy (anonymous):

Plz i need some help to fill in the blank spaces

OpenStudy (anonymous):

Sin² a = ( 1 - cos2a ) / 2

OpenStudy (anonymous):

is it sin(4x)^2 or sin(4x)^4?

OpenStudy (anonymous):

sin(4x)^4 My appologise

OpenStudy (anonymous):

Chlorophyll........so how do i use this formula u presented to solve the question

OpenStudy (anonymous):

Sin²(4x) or sin^4 (4x) ?

OpenStudy (anonymous):

(sin(4x))^4

OpenStudy (anonymous):

I still have the trouble to understand the question, but I'm thinking: Use the half angle formula to solve (Sin(4x))^2=________-1/2cos(_________x)+1/8cos(16x)

OpenStudy (anonymous):

This is going to require a couple applications of these two identities: \[\sin(2 \theta)=2\sin(\theta)\cos(\theta)\] \[\sin^2(\theta)=\frac{1}{2}(1-\cos(\theta))\]

OpenStudy (anonymous):

(Sin(4x))^4=________-1/2cos(_________x)+1/8cos(16x)

OpenStudy (anonymous):

its raise to the power of 4 not 2..........plz

OpenStudy (anonymous):

Yes, that's why you have to use those identities twice.

OpenStudy (anonymous):

Though there are also what are called power-reducing identities, which might be more direct. Either way it's going to get a little messy. Gimme a second to find the power-reducing identity for sine..

OpenStudy (anonymous):

OK, this will probably line up more directly with what you're doing: \[\sin^4(\theta)=\frac{3-4\cos(2 \theta)+\cos(4 \theta)}{8}\] Just replace Θ with 4x.

OpenStudy (anonymous):

@CliffSedge I have no idea about the existence of formula, but it works nicely :)

OpenStudy (anonymous):

Yeah, I usually forget they exist and have to derive them fresh each time from the few I can remember. It's a pain in the neck.

OpenStudy (anonymous):

it sure is

OpenStudy (anonymous):

put i appreciate the help

OpenStudy (anonymous):

n.p. Maybe some other time you can try deriving that power-reducing formula from the half-angle and double-angle identities just to exercise your algebra. ;-)

OpenStudy (anonymous):

Yeah MAYBE